# Intensified optical camera with Timepix4 readout

### v3

@@ 2 Methods
For the tests described here, we developed an optical camera based on the Timepix4 chip bump-bonded to a silicon sensor that is adapted for optical photons. This sensor is identical to the one used in optical cameras based on Timepix3. A housing with a C-mount flange ensures that an intensifier or lenses can be correctly mounted and focused.
@@ 2.2 Optical Timepix cameras
In the optical Timepix4 camera used for these experiments, a Timepix3 optical sensor, which is smaller in area than the full Timepix4 chip, was bump-bonded to Timepix4 and read out with SPIDR4 electronics. The leakage current of the sensor was larger than usual, which is likely caused by the edges of the sensor touching parts of the Timepix4 solder bump array. We therefore kept the bias voltage limited to 70 V in the experiments, while typically it can be safely increased to 100 V. The camera had mounting provisions for the image intensifier packaged as Photonis CricketTM, which integrates the intensifier, power supply, and back-end optics between the intensifier output and optical sensor. The camera also implements up to four time-digital-converters (TDC) employing some of the peripheral Timepix4 pixels for precise time-stamping of external signals. We employed it to time-stamp the pulse generator synchronization signal used to trigger the 90 ps laser flash.
@@ 2.3 Optical setup
The 450 nm laser (Thorlabs GSL45A) was triggered with a pulse generator at 185 kHz. The laser pulse had the minimal allowed pulse duration of 90 ps. The produced light flash had provisions for attenuation before collimation onto the camera. In practice, the attenuation was used only in configuration with the intensifier. The right half of Figure 1 shows the camera (x,y) occupancy map with the intensifier for a well-focused beam.
@@ 3.1 Non-intensified camera
First we performed tests of the optical Timepix4 camera without the intensifier. In this configuration the camera is not single-photon sensitive, but detection can be achieved by a sufficiently bright flash. After adjusting the focusing a small spot corresponding to several pixels was illuminated on the sensor. We selected a single pixel in this spot, and all results presented below were obtained for this single pixel, so we did not need to account for the pixel-to-pixel variability issues. We assumed and relied on the previous experience that the pixel behaviour is uniform enough across the chip.
@@ 3.2 Intensified camera
Figure 4: Left: Distribution of ToT versus time difference between the laser and pixel. The TOT selection in a 200 ns range, from 2000 to 2200 ns, is indicated with red dashed lines. Right: Distribution of time differences between the laser synchronization signal and ToA of one of the hit Timepix4 pixels with TOT selection in a 200 ns range fit with a Gaussian function. The measurements were performed at a bias voltage of 70 V.
@@ 3.3 Timewalk correction
We performed timewalk correction by fitting the time difference distributions with a Gaussian function in 92.5 ns slices of ToT. Figure 5 shows the results of the fit for the mean values as a function of ToT in the left part of the figure. The mean values are marked with red dots on the top of the time difference versus the ToT distribution. The right part of the figure shows the two-dimensional distribution of time difference versus ToT after the timewalk correction. The measurements were performed at a bias voltage of 70 V. One can see that the timing jitter is larger for smaller values of ToT, even after the timewalk correction.
@@ 3.5 Dependence on bias voltage
Photons emitted by the P47 scintillator in the intensifier are converted to photoelectrons near the sensor surface. Charge collection from the surface depends on the depletion of silicon in this area and, therefore, depends on the applied bias voltage. We measured the dependence of the collection time behaviour on the bias voltage by acquiring datasets for six different voltages: 20, 30, 40, 50, 60, and 70 V in the configuration with an intensifier. Figure 7 shows the ToT versus time difference distributions and individual time difference distributions for all values of the bias voltage. We see that the measured time difference decreases from over 100 ns values for a bias of 20 V when the sensor should start to be fully depleted to the values around 50 ns for a maximum bias voltage of 70 V. The latter delay includes the drift time through 300 micron of depleted silicon and additional time offset between the signals. In all cases the timewalk effect is clearly visible. The time difference distributions in the right part of Figure 7 were selected with ToT in the 2000 - 2200 ns range indicated with red lines in the left part of the figure.
@@
Figure 8: Left: Time delay between the laser synchronization signal and ToA of one of the hit Timepix4 pixels as a function of bias voltage. The pixel is activated with a flash of light from the laser with TOT selection in a 200 ns slice. Right: Measured time resolution with TOT selection in a 200 ns slice as a function of bias voltage. The 2000-2200 ns ToT slice is indicated with red lines in Figure 7.
An interesting observation is also that the time resolution exhibits a dependence on the applied bias voltage even for identical ToT values. This behavior suggests that the sensor-related effects, such as, for example, increased time jitter arising from charge diffusion during carrier drift in the depleted silicon region might play a role at the lower bias voltages.
@@ 4 Conclusions
Table 1 presents a comparison of timing resolutions (RMS) achieved in this study for different configurations of the Timepix4 camera. The resolution deteriorates after the addition of the intensifier, mostly due to the appearance of a scintillator as an intermediate step in the amplification chain. The P47 scintillator typically used in fast image intensifiers has a rise time of about 7 ns and, therefore, can limit the time resolution in situations with insufficient gain. Faster scintillators are becoming available [53, 57], which can be used in intensifiers and also for similar MCP-based amplification schemes employed for ion and electron detection. Another opportunity to increase the signal amplitude with the aim of improving time resolution is to employ optical sensors based on low-gain avalanche diodes (LGAD), which typically provide a gain of 5 to 10 [58, 59, 60]. Work is in progress to implement and evaluate these options.
@@ Acknowledgments
This research was supported by the Czech Science Foundation (GACR) under Project No. 25-15534M and Czech Ministry of Education, Youth and Sports Project No. LM2023040 CERN-CZ. We thank Lou-Ann Pestana De Sousa for help with preparation of manuscript.

### v4

@@ 2 Methods
For the tests described here, we developed an optical camera based on the Timepix4 chip bump-bonded to a silicon sensor that is adapted for optical photons. This sensor is identical to the one used in optical cameras based on Timepix3. A housing with a C-mount flange ensures that an intensifier or lenses can be correctly mounted and focused on to the sensor.
@@ 2.2 Optical Timepix cameras
In the optical Timepix4 camera used for these experiments, a Timepix3 optical sensor, which is smaller in area than the full Timepix4 chip, was bump-bonded to Timepix4 and read out with SPIDR4 electronics. The leakage current of the sensor was larger than usual, which is likely caused by the edges of the sensor touching parts of the Timepix4 solder bump array. We therefore kept the bias voltage limited to 70 V in the experiments, while typically it can be safely increased to 100 V. The camera had mounting provisions for the image intensifier packaged as Photonis CricketTM, which integrates the intensifier, power supply, and back-end optics between the intensifier output and optical sensor. The camera also implements up to four time-to-digital-converters (TDC) employing some of the peripheral Timepix4 pixels for precise time-stamping of external signals. We employed it to time-stamp the pulse generator synchronization signal used to trigger the 90 ps laser flash.
@@ 2.3 Optical setup
The 450 nm laser (Thorlabs GSL45A) was triggered with a pulse generator at 185 kHz. The laser pulse had the minimal allowed pulse duration of 90 ps (FWHM). The produced light flash had provisions for attenuation before collimation onto the camera. In practice, the attenuation was used only in configuration with the intensifier. The right half of Figure 1 shows the camera (x,y) occupancy map with the intensifier for a well-focused beam.
@@ 3.1 Non-intensified camera
First we performed tests of the optical Timepix4 camera without the intensifier. In this configuration the camera is not single-photon sensitive, but detection can be achieved by a sufficiently fast and bright flash. After adjusting the focusing a small spot corresponding to several pixels was illuminated on the sensor. We selected a single pixel in this spot, and all results presented below were obtained for this single pixel, so we did not need to account for the pixel-to-pixel variability issues. We assumed and relied on the previous experience that the pixel behaviour is uniform enough across the chip.
@@ 3.2 Intensified camera
Figure 4: Left: Distribution of ToT versus time difference between the laser and pixel. The TOT selection in a 200 ns range, from 2000 to 2200 ns, is indicated with red dashed lines. Right: Distribution of time differences between the laser synchronization signal and ToA of one of the hit Timepix4 pixels with TOT selection in the 200 ns range fit with a Gaussian function. The measurements were performed at a bias voltage of 70 V.
@@ 3.3 Timewalk correction
We performed timewalk correction by fitting the time difference distributions with a Gaussian function in 92.5 ns slices of ToT. Figure 5 shows the results of the fit for the mean values as a function of ToT in the left part of the figure. The mean values are marked with red dots on the top of the time difference versus the ToT distribution. The right part of the figure shows the two-dimensional distribution of time difference versus ToT after the timewalk correction. The measurements were performed at a bias voltage of 70 V. One can see that the timing jitter is larger for smaller values of ToT.
@@ 3.5 Dependence on bias voltage
Photons emitted by the P47 scintillator in the intensifier are converted to photoelectrons near the sensor surface. Charge collection from the surface depends on the depletion of silicon in this area and, therefore, depends on the applied bias voltage. We measured the dependence of the collection time behaviour on the bias voltage by acquiring datasets for six different voltages: 20, 30, 40, 50, 60, and 70 V in the configuration with an intensifier. Figure 7 shows the ToT versus time difference distributions and individual time difference distributions for all values of the bias voltage. We see that the measured time difference decreases from over 100 ns values for a bias of 20 V when the sensor should start to be depleted to the values around 50 ns for a maximum bias voltage of 70 V. The latter delay includes the drift time through 300 micron of depleted silicon and additional time offset between the signals. In all cases the timewalk effect is clearly visible. The time difference distributions in the right part of Figure 7 were selected with ToT in the 2000 - 2200 ns range indicated with red lines in the left part of the figure.
@@
with $d$ the sensor thickness, $t_{0}$ the constant offset due to other effects and $\mu_{h}$ the hole mobility. The dependance of equation 3.1 is only valid for a fully depleted sensor in the low-field region. Based on the sensor properties, full depletion is expected around 27 V. The sensor thus will not be fully depleted for the lowest applied bias of 20 V. In this case, the charge moves through the undepleted region by diffusion rather than drift, which is a slower process and thus introduces additional delays. A precise calculation of this requires detailed knowledge of the sensor physics and is outside of the scope of this work. We therefore omit this measurement from the following analysis.
@@
Figure 8: Left: Time delay between the laser synchronization signal and ToA of one of the hit Timepix4 pixels as a function of bias voltage. The pixel is activated with a flash of light from the laser with TOT selection in a 200 ns slice. The delay values are fit with a theoretical curve, see the text. The point at 20 V was not used in the fit. Right: Measured time resolution with TOT selection in a 200 ns slice as a function of bias voltage. The 2000-2200 ns ToT slice is indicated with red lines in Figure 7.
The observed time-walk behaviour is primarily governed by the pixel preamplifier peaking time, with an additional contribution from the finite rise and decay time of the P47 scintillator in the intensifier, all convoluted with the discriminator response. While drifting carriers in the depleted silicon introduce some voltage-dependent variation, the overall shape of the time-walk curve remains largely unchanged across the relevant bias range, indicating that weighting-field effects and carrier mobilities play a secondary role. In configurations using n-on-p sensors with electron collection, the intrinsic jitter could in principle be reduced for large signals; however, in the intensified optical scheme the scintillator emission time is the dominant limitation, so only a modest improvement would be expected. Properties of Timepix4 front-end electronics and its implication on time measurement are discussed in detail in [50, 57, 58].
@@ 4 Conclusions
Table 1 presents a comparison of timing resolutions (RMS) achieved in this study for different configurations of the Timepix4 camera. The resolution deteriorates after the addition of the intensifier, mostly due to the appearance of a scintillator as an intermediate step in the amplification chain. The P47 scintillator typically used in fast image intensifiers has a rise time of about 7 ns and, therefore, can limit the time resolution in situations with insufficient gain. Faster scintillators are becoming available [53, 59], which can be used in intensifiers and also for similar MCP-based amplification schemes employed for ion and electron imaging. Another opportunity to increase the signal amplitude with the aim of improving time resolution is to employ optical sensors based on low-gain avalanche diodes (LGAD), which typically provide a gain of 5 to 10 [60, 61, 62]. Work is in progress to implement and evaluate these options.
@@ Acknowledgments
This research was supported by the Czech Science Foundation (GACR) under Project No. 25-15534M and Czech Ministry of Education, Youth and Sports Project No. LM2023040 CERN-CZ. This work was supported by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL Programme. We thank Lou-Ann Pestana De Sousa for help with preparation of the manuscript.